OpenStudy (anonymous):

Given f(x) = the quantity of x minus 7, divided by 3, solve for f-1(4).

OpenStudy (anonymous):

-1 19 33 37

OpenStudy (anonymous):

@magbak

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\[\large f(x) = \dfrac{x-7}{3}\] like this ?

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

you want to find \(\large f^{-1}(4)\) that means you need to find what value of x gives you 4 : \[\large 4 = \dfrac{x-7}{3}\] solve x

OpenStudy (anonymous):

how do I do that

ganeshie8 (ganeshie8):

cross multiply, you get : \[\large 12 = x-7\] can u take it from here ? :)

OpenStudy (magbak):

the answer is 19